Need to solve: 1/2(2k-4)=3(k+2)-3k
Answer by checkley77(12844)

1/2(2k-4)=3(k+2)-3k
K-2=3K+6-3K
K-2=6
X=6+2
K=8
PROOF:
1/2(2*8-4)=3(8+2)-3*8
1/2(16-4)=3*10-24
1/2*12=30-24
6=6